OpenStudy (anonymous):

Choose the correct slope of the line that passes through the points (1, −3) and (3, −5)

OpenStudy (anonymous):

(1, -3) (3, -5) -3-(-5)= 2/-2 = -1 so youre slope is -1 1- 3 to find the y intercept: y=mx+b -3=-1(1)+b -3=-1+b -2=b so your y-intercept is -2 re-write the equation: y=-1x-2 there you go :* hope that helps